OpenStudy (anonymous):

a bag contains 3 red 4 white and 5 green balls. Three balls are selected without replacement. Find the probability that the three balls chosen are: (a) one of each color If the selection of the three balls was carried out 1100 times, how often would you expect to choose: (b) 3 red balls (e) one of each color

OpenStudy (anonymous):

First of all: The number of ways to select 3 balls is: \[ ^{3+4+5}P_3=^{12}P_3=\frac{12!}{(12-3)!} = \frac{12!}{9!} \]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

(a) The number of ways to select one of each color, where order doesn't matter, is: \[ 3\times 4\times 5 \]Factoring in order (\(3!\) permutations) we get: \[ 3!\times (3\times 4\times 5) \]

OpenStudy (anonymous):

do u mean 1320

OpenStudy (anonymous):

The probability for one of each color is: \[ \frac{3!\times 3\times 4\times 5}{\frac{12!}{9!}} \]

OpenStudy (anonymous):

For 3 red balls, the total number of ways is: \[ 3\times 2\times 1 \]There is only one possible permutation, so this alone is enough.

OpenStudy (anonymous):

Thus the probability of 3 red balls is: \[ \frac{3\times 2\times 1}{\frac{12!}{9!}} \]

OpenStudy (anonymous):

yes ur correct

OpenStudy (anonymous):

The expected number of times that something happens \(n\) times with probability \(p\) is:\[ n\times p \] In this case \(n = 1100\).

OpenStudy (anonymous):

Where is (c) and (d)?

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

How did you know I was correct?

OpenStudy (anonymous):

i have got answers at the back of my textbook but there wasn't any explanation